Given that (z_1=r_1(cos∅_1+isin∅_1 and (z_2=r_2(cos∅_2+isin∅_2) evaluate (z_1z_2)
1
Expert's answer
2019-10-22T05:58:56-0400

"z_1*z_2 = r_ \n1\n (cos \u03b8_ \n1\n + j sin \u03b8_ \n1\n ) \\displaystyle\\times{r}_{{2}}{\\left( \\cos{\\ }\\theta_{{2}}+{j}\\ \\sin{\\ }\\theta_{{2}}\\right)}\n )\n= r_ \n1\n r_ \n2\n\u200b\t\n (cos[\u03b8_ \n1\n\u200b\t\n +\u03b8_ \n2\t\n ] \\displaystyle{\\left.+{j}\\ \\sin{{\\left[\\theta_{{1}}+\\theta_{{2}}\\right]}}\\right)}\n ])" is the final answer.